OpenStudy (anonymous):

The sum of the first and second of three integers is 154. Find the three even integers.

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

I'm guessing integer 2 = (integer 1)+2, since they are all even?

OpenStudy (anonymous):

umm... i guess that is a reasonable assumption yes.

OpenStudy (anonymous):

*even and (mostly) consecutive

OpenStudy (anonymous):

\[I_1 + I_2 = 154\] \[I_2 = I_1+2\] \[I_3 = I_2 +2\]

OpenStudy (anonymous):

so that would mean that i 2 is 156?

OpenStudy (welshfella):

first even number = n n + n + 2 = 154 2n = 152

OpenStudy (anonymous):

79 is i1?

OpenStudy (welshfella):

no n = 152/2

OpenStudy (welshfella):

i1 = 76

OpenStudy (anonymous):

so if n=76 that would mean I2 is 78

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

and that would mean i3 is,,, 80?

OpenStudy (welshfella):

right

OpenStudy (anonymous):

and that would be the answers?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

ok! thanks again!

OpenStudy (welshfella):

yw